# Language-free Compositional Action Generation via Decoupling Refinement

### v1

@@ Abstract
Composing simple elements into complex concepts is crucial yet challenging, especially for 3D action generation. Existing methods largely rely on extensive neural language annotations to discern composable latent semantics, a process that is often costly and labor-intensive. In this study, we introduce a novel framework to generate compositional actions without reliance on language auxiliaries. Our approach consists of three main components: Action Coupling, Conditional Action Generation, and Decoupling Refinement. Action Coupling utilizes an energy model to extract the attention masks of each sub-action, subsequently integrating two actions using these attentions to generate pseudo-training examples. Then, we employ a conditional generative model, CVAE, to learn a latent space, facilitating the diverse generation. Finally, we propose Decoupling Refinement, which leverages a self-supervised pre-trained model MAE to ensure semantic consistency between the sub-actions and compositional actions. This refinement process involves rendering generated 3D actions into 2D space, decoupling these images into two sub-segments, using the MAE model to restore the complete image from sub-segments, and constraining the recovered images to match images rendered from raw sub-actions. Due to the lack of existing datasets containing both sub-actions and compositional actions, we created two new datasets, named HumanAct-C and UESTC-C, and present a corresponding evaluation metric. Both qualitative and quantitative assessments are conducted to show our efficacy. [^1] Our code can be found at: https://github.com/XLiu443/Language-free-Compositional-Action-Generation-via-Decoupling-Refinement
Figure 1: The comparison between original separate action generation and compositional action generation. The compositional method aims to amalgamate two distinct sub-actions, like “walking” and “drinking”, into a singular, unseen composite concept—illustrated here as “walking+drinking”. This study emphasizes simultaneous composition, exemplified by the integration of concurrent actions such as “walking while drinking”.
@@ 1 Introduction
Humans possess the remarkable ability to generate new knowledge by synthesizing it from prior experiences [16]. This capacity for compositional generation is not only a fundamental hallmark of intelligence but also plays an indispensable role in the pursuit of general artificial intelligence. This paper presents our efforts directed towards facilitating these compositional generation capabilities specifically in the sphere of 3D action generation. Although substantial strides have been made in synthesizing realistic 3D motion sequences [23, 9, 1], the task of generating complex 3D actions from basic elements presents its own unique challenges. This complexity arises from the inherent difficulty in understanding the elemental latent concepts and the principles of compositions without any form of supervision.
@@
Recent developments [30, 38, 39] generate actions conditioned by text to reconcile the semantic space of language with the space of motion. For instance, MotionCLIP [30] operates by positioning semantically analogous motions within the CLIP [25] latent space into close proximity. Similarly, MotionDiffuse [38] deploys CLIP for partial model initialization and progresses to learn a text-conditional diffusion model utilizing data annotated with language. However, these methods have notable limitations. Firstly, they depend on an exhaustive collection of composite actions and corresponding high-quality text annotations, a process that is not only costly but also labor-intensive. Secondly, the composite actions gathered in this manner are predominantly temporally structured, like "walk and then run." As a result, they fall short in generating simultaneous compositional actions such as "walking while drinking."
@@
In this paper, we consider a scenario that is more challenging yet more commonplace, where neither compositional action data nor annotated textual descriptions are provided and simultaneous compositions are involved. To surmount these obstacles, we propose a new framework aimed at learning the compositional action generation model using sub-actions solely. This framework is constituted of three principle components: Action Coupling, Conditional Action Generation, and Decoupling Refinement.
@@
  • Action Coupling is motivated by our observation that humans often concurrently perform two actions using separate body parts, such as “drinking while walking” in Figure 1. It is unreasonable for one body part to simultaneously execute two actions, such as the impossible combinations “calling and singing” and “drinking and throwing”. Thus, we propose implementing an energy model to identify the most active body parts during sub-actions, and subsequently, employing attention masks to couple two actions that involve different active body parts. These coupled actions can serve as the pseudo-compositional data for the model training.
@@
  • Conditional Action Generation is predicated on a conditional generative model, which uses a pair of sub-action labels as its conditioning variables. In particular, we utilize a Conditional Variational Autoencoder (CVAE) [29] model to develop a latent space. From this space, we can sample a devise array of compositional actions, given a pair of sub-action labels.
@@
  • Decoupling Refinement is introduced to improve the quality of generated actions, aiming to ensure semantic consistency between sub-actions and compositional actions. This is particularly crucial given that training data may become distorted during the coupling process. Specifically, we project the generated 3D actions into a 2D space and then separate them into two sub-segments using attention masks. We then employ a self-supervised pre-trained model Mask AutoEncoder (MAE) [10] to reconstruct the complete image from sub-segments. Finally, we impose constraints on these reconstructed images to ensure they align with images rendered from the original sub-actions.
@@
Owing to the absence of existing datasets that comprise both sub-actions and compositional actions, we have devised two new datasets for effective benchmarking. These newly constructed datasets are based on the HumanAct12 [9] and UESTC [14] datasets and are thus named HumanAct-C and UESTC-C. In these datasets, the training sets only consist of sub-actions, while the testing sets are assembled by manually applying an attention mask to each pair of sub-actions to compose them. A subsequent manual filtering process ensures the quality of these composed action pairs in the testing sets. In addition, we introduce a novel evaluation metric inspired by the typical Frechet Inception Distance (FID) metric. We utilize the manually selected testing samples to train an action recognition model Spatio-Temporal Graph Convolutional Networks (STGCN) [34]. For the generated actions, we employ this model as a classifier and compute the FID score. A comprehensive series of both quantitative and qualitative evaluations on two datasets substantiate the effectiveness of our proposed methodology in addressing the challenges of compositional action sequence generation.
Contributions. (1) We introduce a novel task aimed at generating simultaneous compositional 3D action sequences using substantially fewer annotations. (2) We put forward a unique framework that integrates an energy model-based action coupling technique and a semantic-aligned decoupling refinement process. (3) We have constructed two new datasets and offer a new evaluation metric to facilitate benchmarking this emerging task. (4) Finally, through extensive quantitative and qualitative evaluations, we demonstrate the effectiveness of our methodology.
@@ 2 Related Work
Human action synthesis. Human action synthesis, a pivotal task in video understanding, has potential applications in fields requiring virtual or robotic characters. The primary methodologies employed in this domain can be categorized into generative adversarial network (GAN) based methods [7], flow-based methods [28], and variational autoencoder (VAE) frameworks [15]. Notably, GAN-based techniques, such as BAN [40] and Kinetic-GAN [5], combine Bayesian inference with GAN or use graph convolutional networks (GCN) to model the spatial-temporal structure of human motion, respectively. Normalizing flows [36, 11] are also frequently used in human motion generation. For the CVAE [29] based generative framework, Motion2Action applies a GRU-based architecture to model action, while ACTOR further proposes a Transformer-based architecture.
Compositional Image Generation. In the domain of image generation, there have been some developments to equip models with compositional capability. One prevalent approach involves learning disentangled factors of variation [31, 13, 6, 21, 12, 17, 33, 2]. This method decomposes the observed object into distinct latent factors and then recombines these factors to generate new concepts. Another method focuses on learning a disentangled latent space in domain translation [3, 24], although this requires model re-training for different compositions. Some methods involve learning pixel-level attention masks for each factor, followed by assembly in a new layout [8]. Besides, some approaches employ large-scale language pre-trained models [4, 25, 27, 26] to achieve compositional image generation, but they often require a massive amount of data, running into the millions or even billions. Unlike these methods, our work targets a more complex task—generating simultaneous compositional actions using minimal annotations, a paradigm that current methods don’t address effectively.
@@ 3 Framework
Figure 2: The overall pipeline of our framework. It contains three main components, including Action Coupling (in orange), Conditional Action Generation(in green), and Decoupling Refinement (in blue). These processes involve identifying active regions of sub-action sequences using motion energy, which serves as an attention mask. These masked sub-actions and labels are mixed to create pseudo-compositional actions for training a Transformer-based conditional generation model. Generated compositional actions are then converted into 2D images and decoupled into masked images. A pre-trained MAE model is used to recover these images, ensuring the generated compositional actions preserve their semantic information.
@@ 3.2 Action Coupling
Some insights: Before describing the functions $f_{x},f_{y}$, we introduce some insights on human actions, which will motivate our following designs. First, one body part cannot simultaneously execute two actions. For example, it is reasonable to perform “drinking while walking” but unreasonable for “calling and singing”. It means that the simultaneous compositional actions would include two body parts to perform different sub-actions, which motivates us to build two attention masks to focus on different body parts of sub-actions when coupling.
@@
$\vspace{-5mm}\mathbf{E}_{n}=\frac{1}{|\Omega|(T-1)}\sum_{\mathbf{y}_{n}\in\Omega}\sum_{t=2}^{T}||\mathbf{y}_{n,t}-\mathbf{y}_{n,t-1}||^{2},\vspace{-5mm}$ (2)
$\vspace{-5mm}\tilde{\mathbf{y}}_{n}$ $=\frac{\lambda\textbf{E}^{i}_{n}\mathbf{y}^{i}_{n}}{\lambda\textbf{E}^{i}_{n}+(1-\lambda)\textbf{E}^{j}_{n}}+\frac{(1-\lambda)\textbf{E}^{j}_{n}\mathbf{y}^{j}_{n}}{\lambda\textbf{E}^{i}_{n}+(1-\lambda)\textbf{E}^{j}_{n}}.$ (3)
@@ 3.3 Conditional Action Generation
$-\lambda\mathbb{E}_{q_{\phi}(\mathbf{z}|\tilde{\mathbf{x}},\tilde{\mathbf{y}})}[\log g_{\psi}(\tilde{\mathbf{y}}|\tilde{\mathbf{x}},\mathbf{z})].\vspace{-5mm}$
@@ 3.4 Decoupling Refinement
$\vspace{-0.2cm}\mathcal{A}^{i}(\textbf{pix})=\sum_{n=1}^{N}\frac{\textbf{E}^{i}_{n}}{||\textbf{pix}-\textbf{pix}^{i}_{n}||^{2}},\vspace{-0.2cm}$ (6)
@@
Table 1: Comparison with baseline methods on UESTC-C and HumanAct-C. Lower FID-C is better. Acc denotes the action recognition accuracy (higher is better). Div and Mul respectively denote the overall and per-class diversity.

### v2

@@ Abstract
Composing simple actions into complex actions is crucial yet challenging. Existing methods largely rely on language annotations to discern composable latent semantics, which is costly and labor-intensive. In this study, we introduce a novel framework to generate compositional actions without language auxiliaries. Our approach consists of three components: Action Coupling, Conditional Action Generation, and Decoupling Refinement. Action Coupling integrates two sub-actions to generate pseudo-training examples. Then, a conditional generative model, CVAE is employed to facilitate the diverse generation. Decoupling Refinement leverages a self-supervised pre-trained model MAE to ensure semantic consistency between sub-actions and compositional actions. Due to the lack of existing datasets containing both sub-actions and compositional actions, we create two new datasets, named HumanAct-C and UESTC-C. Both qualitative and quantitative assessments are conducted to show our efficacy.  [^1] Our code can be found at: https://github.com/XLiu443/Language-free-Compositional-Action-Generation-via-Decoupling-Refinement
11footnotetext: Corresponding author.
Index Terms—  compositional generation, 3D action generation
@@ 1 Introduction
Humans possess the remarkable ability to generate new knowledge from prior experiences [1]. This capacity for compositional generation plays an indispensable role in artificial intelligence. This paper presents our efforts towards facilitating these compositional generation capabilities in 3D action generation. Although substantial strides have been made in synthesizing realistic 3D motion sequences [2, 3, 4], generating complex 3D actions from basic elements presents its own unique challenges. This complexity arises from the difficulty in understanding the elemental latent concepts and the principles of compositions without any forms of supervision.
Fig. 1: The comparison between separate action generation and compositional action generation. The compositional method aims to amalgamate two sub-actions, like “walking” and “drinking”, into a simultaneous, unseen composite concept - “walking + drinking”.
@@
Recent developments [5, 6, 7] generate actions conditioned by text to reconcile the semantic space of language with the space of motion. MotionCLIP [5] operates by positioning semantically analogous motions within the CLIP [8] latent space into close proximity. MotionDiffuse [6] deploys CLIP for partial model initialization and progresses to learn a text-conditional diffusion model utilizing language annotations. However, these methods have notable limitations. Firstly, they depend on an exhaustive collection of composite actions and corresponding high-quality text annotations, a process that is not only costly but also labor-intensive. Secondly, the composite actions gathered in this manner are predominantly temporally structured, like "walk and then run." As a result, they fall short in generating simultaneous compositional actions such as "walking while drinking."
@@
In this paper, we consider a scenario that is challenging yet commonplace, where neither compositional action data nor annotated textual descriptions are provided and simultaneous compositions are involved. To surmount these obstacles, we propose a new framework aimed at learning the compositional action generation model using sub-actions solely. This framework is constituted of three components: Action Coupling, Conditional Action Generation, and Decoupling Refinement.
@@
  • Action Coupling is motivated by our observation that humans often concurrently perform two actions using separate body parts, such as “drinking while walking” in Figure 1. Thus, we propose to couple two sub-actions that involve different active body parts serving as the pseudo-compositional actions.
@@
  • Conditional Action Generation is predicated on a conditional generative model, which uses a pair of sub-action labels as its conditioning variables. In particular, we utilize a Conditional Variational Autoencoder (CVAE) [9] model to develop a latent space for motion generation.
@@
  • Decoupling Refinement is introduced to improve the quality of generated actions, aiming to ensure semantic consistency between sub-actions and compositional actions. This is particularly crucial given that training data may become distorted during the coupling process. Specifically, we render and split the generated 3D skeletons into two sub-segments and use a pre-trained MAE model to recover the missing sections.
@@
Owing to the absence of existing datasets that comprise both sub-actions and compositional actions, we devise two new datasets as benchmarks constructed on the HumanAct12 [3] and UESTC [10] datasets. In these datasets, the training sets only consist of sub-actions, while the testing sets are assembled by manually applying an attention mask to each pair of sub-actions to compose them. Comprehensive quantitative and qualitative evaluations on two datasets substantiate the effectiveness of our proposed methodology in addressing the challenges of compositional action sequence generation.
@@ 2 Related Work
Human Motion Generation Human motion generation has potential applications in fields requiring virtual or robotic characters. The primary methods employed in this domain can be categorized into generative adversarial network (GAN) based methods [11], flow-based methods [12], and variational autoencoder (VAE) frameworks [13]. Notably, BAN [14] and Kinetic-GAN [15] apply GAN to model human motions while Action2motion [3] and ACTOR [2] propose different VAE-based architectures. Recent methods [5, 6, 7] have proposed using neural language as the conditional factor, which leverages the semantic structure inherent in language to empower generation ability. Unlike these methods, our work targets a complex task—generating simultaneous compositional actions using minimal annotations, a paradigm that current methods don’t address effectively.
Compositional Image Generation. In the field of image generation, there have been some developments to equip models with compositional capability. One prevalent approach involves learning disentangled factors of variation [16, 17]. This method decomposes the observed object into distinct latent factors and then recombines these factors to generate new concepts. Another method focuses on learning a disentangled latent space in domain translation [18, 19]. Other approaches employ large-scale language pre-trained models [8, 20, 21] to achieve compositional image generation, but they often require a massive amount of data, running into the millions or even billions.
@@ 3 Framework
Fig. 2: The pipeline of our framework. It contains three main components, including Action Coupling (in orange), Conditional Action Generation(in green), and Decoupling Refinement (in blue). These processes involve identifying active regions of sub-actions with motion energy, and mixing sub-actions as pseudo-compositional actions for training a conditional generation model. Generated compositional actions are then converted and decoupled into masked images. A pre-trained MAE model is used to recover these images.
@@ 3.1 Problem Definition
We denote a pair of action labels and their corresponding motion sequences as $\mathbf{x},\mathbf{y}\in\mathcal{X}\times\mathcal{Y}$. Here, $\mathbf{x}$ signifies the sub-action label, while $\mathbf{y}=\{\mathbf{y}_{t}\in\mathbb{R}^{N\times 3}|t=1,2,\cdots,T\}$ represents the sequence of action poses. Specifically, each frame $\mathbf{y}_{t}$ comprises a pose that includes N points $\{\textbf{y}_{n,t}|n=1,2,...,N\}$, each of which is a 3D coordinate, with $T$ denoting the sequence length. Beyond this sub-action sequence, there lies an unseen compositional action space $\mathcal{Y}\times\mathcal{Y}\rightarrow\mathcal{\tilde{Y}}$, wherein each composite action is constructed from two sub-actions. The aim is to learn a generative model $g_{\psi}(\mathbf{\tilde{y}}|\mathbf{x}^{i},\mathbf{x}^{j},\mathbf{z})$ that produces a composite action $\mathbf{\tilde{y}}\in\mathcal{\tilde{Y}}$ given two different sub-action labels $\mathbf{x}^{i},\mathbf{x}^{j}$ (here we use the superscript to represent different samples) and a noise vector $\mathbf{z}$.
@@ 3.2 Action Coupling
Action Coupling aims to synthesize compositional action sequences from observed sub-actions. Given two label variables $\mathbf{x}^{i},\mathbf{x}^{j}$ and corresponding sub-action sequences $\mathbf{y}^{i},\mathbf{y}^{j}$, we couple two sub-action sequences to generate pseudo targets in the following procedure:
@@
Mixing rate distribution: In the context of Mix-up methods for data augmentation [22], the parameter $\lambda$ is sampled from $\lambda\sim Beta(\alpha,\alpha)$, where $\alpha\in(0,\inf)$ is a hyper-parameter that regulates the intensity of interpolation. This Beta distribution pushes the $\lambda$ close to 00 or $1$ and away from $0.5$. It is reasonable, as overly mixed images (e.g., where $\lambda=0.5$) may dilute discriminative capacity, thereby contradicting the objective of data augmentation, which is to enhance classification. Conversely, our method is centered around generating composite actions that preserve as much information from sub-actions as possible. In essence, we promote the deep integration of two actions, ideally where $\lambda=0.5$. As such, we employ $\lambda\sim\mathcal{N}(0.5,\delta)$, where $\delta$ denotes the variance of the Gaussian distribution.
Preliminary insights: Prior to exploring the mechanics of $f_{x},f_{y}$, we aim to convey observations about human actions that serve as the motivation for our following designs. Firstly, it is important to recognize that a single body part is incapable of executing two distinct actions simultaneously, which implies that simultaneous compositional actions generally involve two different body parts carrying out different sub-actions. This insight motivates our strategy of crafting two attention masks focusing on disparate body parts during action coupling. Secondly, we find that body parts with the highest motion energy tend to be crucial in defining a particular action. Essentially, significant motion is almost always associated with a body part that plays a key role in a certain sub-action. This insight drives us to calculate motion energy to identify the active body parts.
@@
$\mathbf{E}_{n}=\frac{1}{|\Omega|(T-1)}\sum_{\mathbf{y}_{n}\in\Omega}\sum_{t=2}^{T}||\mathbf{y}_{n,t}-\mathbf{y}_{n,t-1}||^{2},$ (2)
where $\Omega$ denotes a body part, containing a set of joint points $\{\textbf{y}_{n,t}|n\in\Omega\}$, and $n,t$ denotes the $n$th joint point in the $t$th frame. Note that for the joint points in the same body part, the energy values are equal. Then, we define the attention value as the energy value. Finally, we couple two sub-actions with the obtained attention mask $\mathbf{E}_{n}$ and implement functions $f_{y}$ in the second line in Eq (8) as
$\tilde{\mathbf{y}}_{n}$ $=\frac{\lambda\textbf{E}^{i}_{n}\mathbf{y}^{i}_{n}}{\lambda\textbf{E}^{i}_{n}+(1-\lambda)\textbf{E}^{j}_{n}}+\frac{(1-\lambda)\textbf{E}^{j}_{n}\mathbf{y}^{j}_{n}}{\lambda\textbf{E}^{i}_{n}+(1-\lambda)\textbf{E}^{j}_{n}}.$ (3)
@@ 3.3 Conditional Action Generation
This generative process can be learned with CVAE where a variational neural network $q_{\phi}(\mathbf{z}|\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ is introduced to optimize the objective above. The optimization follows the standard variational lower bound:
@@
$-\lambda\mathbb{E}_{q_{\phi}(\mathbf{z}|\tilde{\mathbf{x}},\tilde{\mathbf{y}})}[\log g_{\psi}(\tilde{\mathbf{y}}|\tilde{\mathbf{x}},\mathbf{z})].$
The first term reduces the KL divergence between $q_{\phi}(\mathbf{z}|\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ and $p_{\theta}(\mathbf{z}|\tilde{\mathbf{x}})$, which encourages the consistency of latent variables generated from condition and data. The second term is a reconstruction objective that encourages the model to generate realistic action poses.
@@ 3.4 Decoupling Refinement
Considering the potential compromise in the quality of generated sequences, we introduce Decoupling Refinement, which serves as a constraint to enhance the quality of the generated actions.
3D to 2D rendering: In this study, we utilize the pre-trained MAE [23] as our self-supervised model. However, there is a gap between 3D actions and 2D images (used for MAE). To bridge this gap, we utilize the Skinned Multi-Person Linear (SMPL) body model [24]. Both the original sub-actions and the generated sequences are processed through the SMPL model, producing the dense point clouds representing body surfaces. These 3D outputs are then projected onto a 2D plane using a rendering function. This process can be formulated as $f_{R}:\{\textbf{y}^{i},\textbf{y}^{j},\hat{\textbf{y}}\in\mathbb{R}^{3\times N}\}\rightarrow\{\textbf{v}^{i},\textbf{v}^{j},\hat{\textbf{v}}\in\mathbb{R}^{3\times H\times W}\}$, where v denotes the rendering images with height $H$ and width $W$. To enhance the readability of the resultant 2D images, we standardize the 3D model to ensure rendering is operated from a frontal perspective.
Decoupling: We introduce a reversal of the coupling process, referred as decoupling. which allows us to deconstruct the rendering images of compositional actions back to sub-action components and further facilitate refining generation. Intuitively, for a compositional action, our goal is to preserve the active parts associated with each sub-action. Technically, we propose a 3D-2D attention projection to compute the decoupling attention map using the motion energy $\textbf{E}_{n}$. By utilizing the SMPL projection and the rendering function $f_{R}$, we first transform each 3D joint point $\textbf{y}^{i}_{n}$ into 2D coordinates $\textbf{pix}^{i}_{n}$. The pixels in proximity to active joint points should be prioritized to focus. Hence, we establish a decay rate that is inversely proportional to the square of the pixel distance. We then compute the value for each pixel in the 2D attention map $\mathcal{A}$ by aggregating these decayed attention scores from all $N$ joint points:
$\mathcal{A}^{i}(\textbf{pix})=\sum_{n=1}^{N}\frac{\textbf{E}^{i}_{n}}{||\textbf{pix}-\textbf{pix}^{i}_{n}||^{2}},$ (6)
@@
Constraints: The masked images are fed to a pre-trained MAE model for the inpainting process. Despite the input image being part of a generated composite action, we encourage the model to reconstruct the original sub-actions, denoted as $\hat{\textbf{v}}_{mae}^{i},\hat{\textbf{v}}_{mae}^{j}=f_{mae}(\hat{\textbf{v}}^{i},\hat{\textbf{v}}^{j})$, where $f_{mae}$ denotes the MAE model. It indicates that the attributes of the sub-actions are preserved during the compositional generation process and can be readily disentangled. For all data pairs $\textbf{y}^{i},\textbf{y}^{j}$, the loss function for the decoupling refinement process can be expressed as follows:
@@
Table 1: Comparison with baseline methods on UESTC-C and HumanAct-C. Lower FID is better. Acc denotes the action recognition accuracy (higher is better). Div and Mul respectively denote the overall and per-class diversity.
@@ 4 Experiments
4.1 Dataset & Evaluation Metrics
The HumanAct-C dataset contains a training set with 120 sequences for 12 sub-actions, and a testing set including 9 compositional categories with 900 compositional action sequences. As to the UESTC-C dataset, the testing set comprises 25 compositional categories with $625$ compositional action sequences. The training set is composed of $50$ sequences. Both the training and testing sets consist of the same 10 sub-actions. We employ the Frechet Inception Distance (FID) as our primary evaluation metrics.
4.2 Implementation Details
Network architecture: Our framework is comprised of a Transformer encoder and a decoder, as depicted in Figure 2. The inputs to the Transformer encoder are the pseudo-compositional actions with mixed action embeddings, and the output represents the distribution parameters of the latent space. The Transformer decoder utilizes a combination of the latent vector and the compositional label embedding for its key and value inputs, with a sequence of positional embeddings as the query. And the output is a sequence of 3D actions.
Training details: The loss function of our model comprises three components: the reconstruction loss and the KL loss as per Eq. (12), and the Decoupling Refinement (DR) loss as defined by Eq. (14). The balance rate between these components is set as $21:1e-5:1e-$ 2. We set the frame length to 60, with each frame containing $N=24$ joint points. Both the encoder and the decoder employ eight transformer layers. The AdamW is used as the optimizer with 1e-4 initial learning rate and no learning rate schedule is applied. We trained all models 1500 epochs with batch size 800.
4.3 Quantitative Evaluation
We compare our method with three main baseline methods including Pre-train+Mask [2], Latent Disentanglement [4], and the text-guided method MotionCLIP [5]. Pre-train+Mask is an intuitive compositional generation solution that directly composes two generated actions with attention masks. Table 1 outlines our experimental results, along with those from baselines on the HumanAct-C and UESTC-C datasets. Following ACTOR [2], we report on FID and action recognition accuracy (Acc) as primary evaluation metrics, and also consider overall diversity and per-class diversity as auxiliary diversity metrics. Our method significantly outperforms baseline approaches and text-guided approaches on both datasets. The results underline the potential for our model to generate high-quality compositional motion sequences without extensive text annotations, leading the way for efficient compositional action generation.
4.4 Ablation Studies
We conduct ablation studies to investigate the effectiveness of different components. The performance results are summarized in Table 2. It demonstrates the efficacy of the Decoupling Refinement in boosting the overall performance of our generation model. As revealed by the data in Table 2, the Decoupling Refinement consistently enhances our model’s performance on both the HumanAct-C and UESTC-C datasets. Besides, it can be concluded that the pre-trained text-guided methods cannot be successfully applied to our scenarios. As indicated in Table 1, MotionCLIP [5] struggles to generate compositional actions. In our experiments, we utilize the MotionCLIP model pre-trained on AMASS [25] (we’re unable to train MotionCLIP using our dataset due to the absence of language annotations and compositional data). However, the results show that the model fails to generate satisfactory results. In addition, it has been proven the energy-based attention mask plays a key role in coupling and decoupling. In an experiment where the attention mask was set to $\bm{1}$, indicating equal scores for all body parts, the performance significantly declined. This finding underscores the importance of the attention mask for learning disentangled representations.
Fig. 3: Compositional 3D motion generations with different categories on the HumanAct-C and UESTC-C datasets.
Fig. 4: The qualitative comparisons between our methods and other baseline methods.
4.5 Qualitative Evaluation
In this subsection, we showcase several visualization examples of the generated compositional 3D actions. Figure 3 displays results for four categories from the HumanAct-C and UESTC-C datasets. We see the generation of realistic compositional 3D motions. This validates our method’s efficacy in learning the disentangled latent variable and its applicability to zero-shot compositional action generation. Moreover, Figure 4 compares the generations among the Pre-trained ACTOR [2] + Mask, MotionCLIP [5], and ours.
5 Conclusion
In this paper, we present an innovative framework for generating compositional 3D action sequences, without the need of extensive annotations. We develop two novel datasets, HumanAct-C and UESTC-C, providing valuable resources for future research in this domain. Comprehensive evaluations of these datasets show our methodology’s significant effectiveness in generating compositional actions, opening new avenues for further research in the field of 3D action generation.
Acknowledgment
This work was supported by the following grants:
Shenzhen Key Laboratory of Ubiquitous Data Enabling Grant No. ZDSYS20220527171406015, the UKRI grant of Turing AI Fellowship EP/W002981/1. We would also like to thank the Royal Academy of Engineering and FiveAI.
Appendix for
“Language-free Compositional Action Generation via Decoupling Refinement”
Appendix organization:
[…]
A.1 More Related Work
A.2 More Details of Method
A.2.1 Problem Definition
A.2.2 Action Coupling
$\mathbf{E}_{n}=\frac{1}{|\Omega|(T-1)}\sum_{\mathbf{y}_{n}\in\Omega}\sum_{t=2}^{T}||\mathbf{y}_{n,t}-\mathbf{y}_{n,t-1}||^{2},$ (9)
$\tilde{\mathbf{y}}_{n}$ $=\frac{\lambda\textbf{E}^{i}_{n}\mathbf{y}^{i}_{n}}{\lambda\textbf{E}^{i}_{n}+(1-\lambda)\textbf{E}^{j}_{n}}+\frac{(1-\lambda)\textbf{E}^{j}_{n}\mathbf{y}^{j}_{n}}{\lambda\textbf{E}^{i}_{n}+(1-\lambda)\textbf{E}^{j}_{n}}.$ (10)
A.2.3 Conditional Action Generation
[…]
$-\lambda\mathbb{E}_{q_{\phi}(\mathbf{z}|\tilde{\mathbf{x}},\tilde{\mathbf{y}})}[\log g_{\psi}(\tilde{\mathbf{y}}|\tilde{\mathbf{x}},\mathbf{z})].$
A.2.4 Decoupling Refinement
$\mathcal{A}^{i}(\textbf{pix})=\sum_{n=1}^{N}\frac{\textbf{E}^{i}_{n}}{||\textbf{pix}-\textbf{pix}^{i}_{n}||^{2}},$ (13)
A.3 More Experiments